# How segmented is my network?

### v1

@@ VII Operational Use Cases
Merger and acquisition integration During mergers or acquisitions, organizations must integrate networks while maintaining security boundaries. The segmentedness metric provides a quantitative measure of integration progress. For instance, two networks with pre-merger segmentedness values of $\hat{S_{1}}=0.5$ and $\hat{S_{2}}=0.4$ might temporarily show $\hat{S}_{combined}=0.3$ immediately after connection, reflecting increased connectivity between previously isolated systems. Tracking this metric helps security teams ensure that integration does not inadvertently create an overly flat combined network.
[…]
@@
VIII Practical Considerations for Sampling Node Pairs
@@
Tooling-induced sampling bias In many organizations, sampling node pairs is mediated by security or management tooling, such as firewall policy managers and configuration databases. These tools often expose subsets of the network more readily than others, for example, privileging server-to-server communication over user endpoints or unmanaged devices. Consequently, node pairs may be sampled from the set of entities visible to the tooling rather than uniformly from the underlying network, introducing visibility-driven bias.
Configured vs enforced policy The metric is defined in terms of configured policies, but measurements in practice reflect observed enforcement. The two may differ due to routing constraints, implicit deny rules, or configuration drift.
@@ IX Limitations
Topology independence While structural neutrality is a design goal (Section 3.4), it means the metric does not capture important security-relevant properties such as hub-and-spoke vulnerabilities, critical chokepoints, or segmentation hierarchy.
@@ X Conclusions
We introduce segmentedness as a simple scalar metric for network isolation. By formalizing segmentedness through edge density, we provide practitioners with an interpretable measure of how permissive network communication policies are.
The metric satisfies key properties for practical use: it is normalized for cross-network comparison, scales to networks of arbitrary size, and can be estimated efficiently through randomized sampling with explicit statistical guarantees. Our validation demonstrates that estimates remain unbiased and well-behaved across different network structures.
The main contribution of this paper is in providing a quantitative measure that can be compared across networks and over time. Organizations can now track segmentation trends (e.g., through quarterly measurements requiring fewer than 100 samples), validate security architecture initiatives, and establish empirical baselines for their networks.
Future work includes (a) empirical studies across diverse organizational networks to establish sector-specific benchmarks (e.g., financial services, healthcare, education), (b) extensions incorporating connection constraints and policy attributes beyond binary connectivity, (c) integration with automated network management and security management platforms, and (d) development of complementary structural metrics to provide a comprehensive segmentation assessment.

### v2

@@ VII Operational Use Cases
Merger and acquisition integration During mergers or acquisitions, organizations may be required to integrate networks The segmentedness metric provides a quantitative proxy for network security during the integration process. For instance, two networks with pre-merger segmentedness values of $\hat{S_{1}}=0.5$ and $\hat{S_{2}}=0.4$ might temporarily show $\hat{S}_{combined}=0.3$ immediately after connection, reflecting increased connectivity between previously isolated systems. Tracking this metric helps security teams ensure that integration does not inadvertently create an overly flat combined network.
Zero Trust implementation assessment Organizations implementing zero trust architectures can use segmentedness measurements to validate progress. A traditional network might measure $\hat{S}=0.2$, indicating relatively flat connectivity. As micro-segmentation policies are deployed incrementally as part of a structured zero trust program—first isolating critical assets, then production workloads, then development environments—measurements should show increasing segmentedness (e.g., $\hat{S}=0.3,0.4,0.5$), providing quantitative evidence of architectural improvement aligned with zero trust principles. Further, by quantifying the reduction in global edge density, the segmentedness score provides a direct proxy for the difficulty of unauthorized lateral movement. As the score increases, the number of available reachable paths for an attacker decreases, thereby forcing a reliance on fewer, more heavily monitored chokepoints and significantly raising the effort required to reach critical assets.
@@
A step-by-step procedure for implementing segmentedness measurements in operational environments is provided in Appendix C.
VIII Practical Considerations for Sampling
@@
Tooling-induced sampling bias In many organizations, sampling node pairs is mediated by security or management tooling, such as firewall policy managers and configuration databases. These tools often expose some subsets of the network more readily than others, for example, privileging server-to-server communication over user endpoints or unmanaged devices. Consequently, node pairs may be sampled from the set of entities visible to the tooling rather than uniformly from the underlying network, introducing visibility-driven bias.
Configured vs enforced policy The metric is defined in terms of configured policies, but measurements in practice reflect observed enforcement. The two may differ due to routing constraints, implicit deny rules, or configuration drift from recorded network security policy.
@@
Induced noise A critical operational detail regarding the proposed connectivity test suite is the potential for active probing to trigger security alerts within Intrusion Detection Systems (IDS) [8] or Security Information and Event Management (SIEM) [12] platforms. To mitigate potential disruption to security operations, measurements should ideally be scheduled during maintenance windows or coordinated with security teams to suppress expected alerts. This allows for a controlled environment where the resulting alerts can be monitored and appropriately suppressed from standard escalation paths.
@@ IX Limitations
Topology independence While structural neutrality is a design goal (Section III-D), it means the metric does not capture important security-relevant properties such as hub-and-spoke vulnerabilities, critical chokepoints, or segmentation hierarchy.
@@ X Conclusions
We introduce segmentedness as the first statistically principled scalar metric for network isolation, enabling practitioners to quantify—with explicit confidence intervals—what has previously required only qualitative assessment. By formalizing segmentedness through edge density, we provide practitioners with an interpretable measure of how permissive network communication policies are.
The metric satisfies key properties for practical use: it is normalized for cross-network comparison, scales to networks of arbitrary size, and can be estimated efficiently through randomized sampling with explicit statistical guarantees. Our validation demonstrates that estimates remain unbiased and well-behaved across different networks.
The main contribution of this paper is to provide a quantitative measure that practitioners can deploy immediately without specialized infrastructure or machine learning expertise. Using standard network utilities and basic statistical knowledge, organizations can measure segmentedness with fewer than 100 sampled node pairs and track trends through quarterly measurements, validate security architecture initiatives, and establish empirical baselines—regardless of network size.
@@
Appendix A Graph-based Measures’ Comparison
Table V shows all three graph-based measures from Section II-C for the ten densest networks in the Cisco Secure Workload dataset. The absolute correlation coefficient calculated over this table between Flatness and the Fiedler Value is $0.136$; that between Flatness and Modularity is $0.568$. The low correlation with Fiedler Value ($0.136$) and thr moderate correlation with Modularity ($0.568$) suggests the three measures capture related but distinct aspects of networks (graphs).
TABLE V: Graph-measures comparison: Flatness, Modularity, and Fiedler Value of the 10 densest networks in the Cisco Secure Workload dataset
Graph $F(G)$ Modularity Fiedler Value
g1 0.102 0.077 3.08e-14
g11 0.073 0.380 1.10e-15
g19 0.041 0.358 1.34e-15
g21 0.034 0.310 5.97e-15
g7 0.017 0.629 0.1326
g18 0.012 0.292 3.01e-16
g16 0.011 0.573 0.00e+$0$
g17 0.006 0.627 -4.50e-15
g14 0.005 0.792 1.09e-14
g8 0.002 0.181 -1.07e-13
Appendix B Bayesian Inference for High-segmentedness
The use of the Wald confidence interval in (12) poses a mathematical challenge when the number of observed connections is zero. In such cases, the empirical flatness $\hat{F}$ becomes zero, which causes the standard error to collapse and results in a zero-width confidence interval. This ”zero-width” interval is misleading for security practitioners as it suggests absolute certainty that no communication paths exist, when in reality it is merely a result of the empirical sample and the sample size used.
To address this, we introduce a Bayesian Beta-binomial model to provide a more robust estimation of uncertainty at the boundaries [11]. The Beta distribution is the conjugate prior for the binomial process, making it an ideal choice for updating our beliefs about network flatness based on sampled data.
We incorporate a prior belief that most enterprise networks have an edge density of at least 0.01 (one connection per 100 pairs). This 1% baseline reflects the assumption that even highly segmented enterprise networks maintain minimal operational connectivity (e.g., for monitoring, backup, or administrative access). This prior is modeled using a Beta distribution with hyperparameters $\alpha=1$ and $\beta=99$, which centers our initial expectation at 1%.
Following the empirical observation of $k=0$ connections across $M=97$ sampled pairs, we update these hyperparameters to their posterior values. The posterior distribution becomes $Beta(\alpha^{\prime},\beta^{\prime})$, where the parameters are updated as follows:
$\alpha^{\prime}=\alpha+k=1+0=1$ (16)
$\beta^{\prime}=\beta+M=99+97=196$ (17)
The resulting posterior mean for flatness, representing our new point estimate, is:
$\hat{F}_{Bayes}=\frac{\alpha^{\prime}}{\alpha^{\prime}+\beta^{\prime}}=\frac{1}{197}\approx 0.005$ (18)
Rather than a frequentist confidence interval, we calculate a 95% credible interval using the quantiles of the posterior Beta distribution. For this specific case where $\alpha^{\prime}=1$, the 95% upper bound for flatness ($F_{upper}$) is given by:
$F_{upper}=1-(0.05)^{1/(\alpha^{\prime}+\beta^{\prime}-1)}=1-(0.05)^{1/196}\approx 0.015$ (19)
where the upper bound is derived from the 95th percentile of the posterior $Beta(1,196)$ distribution.
From a security practitioner’s perspective, this means that even with zero observed connections in 97 trials, we can only be 95% certain that the true network flatness is below 0.015. This interpretation provides a more realistic and conservative assessment of risk, acknowledging that rare lateral movement paths may still exist despite not being captured in a small sample. It transforms a mathematically impossible “perfect” score into a statistically defensible claim of high segmentedness.
Appendix C Deployment Guide
Prerequisites
Organizations require three components to measure network segmentedness:
Asset Inventory. A list of networked endpoints with IP addresses, typically maintained through configuration management databases (CMDBs), asset management systems, or network discovery tools.
Connectivity Testing Tools. Standard network utilities for testing communication between node pairs, including ping (ICMP), nc/netcat (TCP/UDP), nmap, or equivalent tools available on Unix-like systems or Windows.
Statistical Computing. Basic capability to compute sample statistics and confidence intervals using Python, R, Excel, or similar tools.
Measurement Procedure
Step 1: Enumerate nodes. Export the list of $n$ network endpoints from the asset inventory system. Each entry should include a unique identifier (IP address or hostname) and be representative of systems subject to segmentation policy.
Step 2: Sample node pairs. Generate $M$ random unordered pairs of distinct nodes. For a 95% confidence interval with $\pm 0.1$ half-width, set $M=97$. Use uniform random sampling without replacement from the set of all possible pairs.
Step 3: Test connectivity. For each sampled pair $(u,v)$, execute the connectivity test suite: ICMP echo request (ping -c 1 -W 1 <target>), TCP connection attempts on common ports (nc -zv -w 1 <target> 80 443), and UDP probes (nc -zuv -w 1 <target> 53). Mark the pair as connected if any test succeeds. Record the number of connected pairs $k$.
Step 4: Compute estimate. Calculate the flatness estimate $\hat{F}=k/M$ and segmentedness estimate $\hat{S}=1-\hat{F}$.
Step 5: Compute confidence interval. The 95% Wald confidence interval is $\hat{S}-0.1,\hat{S}+0.1]$. For the special case where $k=0$, use the Bayesian approach described in Appendix B to avoid a zero-width interval.
Operational Recommendations
Scheduling. Coordinate connectivity testing with security operations teams to suppress expected alerts in IDS and SIEM platforms. Schedule measurements during maintenance windows when possible to minimize operational disruption.
Sampling bias. Ensure the asset inventory is current and complete before sampling. Stale or incomplete inventories may bias measurements toward long-lived or well-managed systems, potentially overestimating segmentedness.
Repeatability. Use the same connectivity test suite across all measurements to ensure comparability over time. Document the specific tests used and any protocol-specific exclusions applied during measurement.
Interpretation. A single measurement establishes a baseline value for network segmentedness. Tracking $\hat{S}$ over time (e.g., quarterly) reveals trends in policy permissiveness and helps identify unintended drift or the impact of architectural changes.